OpenStudy (anonymous):

Find the open interval on which the given power series converges absolutely. E (2^n-2^(-n))(x-1)^n

OpenStudy (amistre64):

take the limit of an/an-1 as n approaches infinity

OpenStudy (amistre64):

\[\lim_{n\to\ inf}\frac{ 2^n(1-2^{-2n})(x-1)^n} {2^{n-1} (1-2^{-2(n-1)})(x-1)^{n-1}}\] \[\lim_{n\to\ inf}\frac{ 2(1-2^{-2n})(x-1)} {1-2^{-2n+1}}\] \[2|x-1|\ \lim_{n\to\ inf}\frac{1-2^{-2n}} {1-2^{-2n+1}}\]

OpenStudy (amistre64):

looks to me like the n parts limit off to 1/1 sooo, if i did this right 2|x-1| < 1 |x-1| < 1/2 -1/2 < x-1 < 1/2 1/2 < x < 3/2